LVMH Moët Hennessy Louis Vuitton SE (LVMH) is a French multinational luxury goods conglomerate headquartered in Paris. The following information is selected from their 2017 annual report.

(€ thousands) 2017 2016
Revenue 42,636 37,600
Operating income 8,113 6,904
Net interest expense and other nonoperating expense 179 432
Net income 5,616 4,363
Operating assets 67,246 58,504
Operating liabilities 26,714 24,340


LVMH has an income tax rate of approximately 30%

a. Compute LVMH’s net operating profit after taxes (NOPAT) for 2017 and 2016.
Round answers to the nearest whole number.

2016
2017


b. Compute LVMH’s net operating profit margin (NOPM) for each year.
Round answers to one decimal place. (Ex: 0.2345 = 23.5%)

2016
2017


c. Compute LVMH’s return on net operating assets (RNOA) for 2017.
Round answers to one decimal place. (Ex: 0.2345 = 23.5%)
Answer #1

Solution:

Part a –

NOPAT means Net Operating Profit after taxes. This profit does not include interest and other non operating expenses but after tax.

i.e. NOPAT = Operating Income x (1 – Tax Rate)

2016 NOPAT = 6,904,000 (1 – 0.30) = 4,832,800

2017 NOPAT = 8,113,000 (1 – 0.30) = 5,679,100

Part b –

Net Operating Profit Margin = NOPAT / Total Revenue x 100

2016 NOPM = 4,832,800 / 37,600,000 * 100 = 12.85% or 12.9%

2017 NOPM = 5,679,100 / 42,636,000 * 100 = 13.32% or 13.3%

Part c –

2017 Return on Net Operating Assets = NOPAT / Average Operating Assets x 100

= 5,679,100 / (67,246,000 + 58,504,000)/2 * 100

= 5,679,100 / 62,875,000 * 100

= 9%

Hope the above calculations, working and explanations are clear to you and help you to understand the concept of question.... please rate my answer...in case any doubt, post a comment and I will try to resolve the doubt ASAP…thank you